The total mass of the hydrogen gas in the Hindenburg zeppelin was 18,000 kg. What volume did the hydrogen occupy? (Assume that the temperature of the hydrogen was 0°C and that it was at a pressure of 1 atm.)
1
Expert's answer
2020-10-29T10:26:24-0400

Given,

Pressure, P = 1 atm

Temperature is 0°C

T = (0 +273)°C

= 273 K


Mass of Hydrogen,W = 18000 kg

= 18000 x 10g

(Since, 1kg = 10g)

Molar mass of Hydrogen, M = 2 g/mole


Number of moles of Hydrogen,

n = (W/M)

Or, n = (18000x103/2) mole

Or, n = 9x10mole


Taking ideal gas behaviour,

We can use Idea gas equation

PV = nRT.

(R= universal gas constant

= 0.082 liter.atm)

Or, V = (nRT/P)

Putting the values,

Or, V =(9x10x0.082 x273/1)

Or, V = 201.474x106


Hence the volume occupied by the given amount of Hydrogen gas = 201.474x10liter